# Spiral of Silence: How is Large Language Model Killing Information Retrieval?—A Case Study on Open Domain Question Answering

### v2

@@ A.7 Attempts to Alleviate the Spiral of Silence
Experimental Setup: To balance the efficiency and effectiveness of the retrieval system, for each set of search results returned by the system, we post-process to acquire the top-5 qualifying documents that are visible to the LLMs. In the source filtering experiment, we employ the Hello-SimpleAI/chatgpt-qa-detector-roberta [^6] https://github.com/Hello-SimpleAI/chatgpt-comparison-detection model to authenticate the origins of the texts within the search results, aiming to retain the first 5 documents identified as human-generated and supply them as input to the LLM’s context. For the content filtering part of the experiment, we apply a selection process based on computing the 3-gram Self-BLEU scores. The specific procedure is as follows: For the top-5 documents returned for each search query, we initially calculate their Self-BLEU scores; if the score exceeds a predetermined threshold (set at 0.4 for this experiment), we then compute the Self-BLEU scores for all possible combinations of 4 documents and select the minimum value among them. This minimum value indicates the maximum individual document contribution to the Self-BLEU score not included in the calculation. Subsequently, we exclude the document contributing the most to the Self-BLEU score and incorporate the next ranked document into the combination, repeating this filtering process until the combination’s Self-BLEU score meets the preset threshold criteria.
@@
Analysis of Experimental Results: From the experimental results, we can observe that: 1) Both approaches yield more stable retrieval outcomes; however, the source filtering method incurs a performance cost. Figure 14 and Figure 15 illustrate the variations in the average retrieval outcomes and QA performance across datasets, before and after the application of two distinct filtering strategies, compared to an unfiltered condition. Observations indicate that by implementing source-based and diversity-based filtering methods, the fluctuation range of the top-5 retrieval results is reduced compared to the non-intervention scenario, suggesting that the filtering mechanisms can bring a more stable retrieval performance for RAG systems. Across the four datasets, the retrieval performance following SELF-BLEU value filtering generally surpasses the unfiltered condition; conversely, the source-based filtering strategy results in an overall performance degradation. This could be attributed to the discriminating model erroneously excluding valid human-generated texts while aiming to eliminate those generated by LLMs. Moreover, in QA tasks, diversity filtering either enhances or maintains QA performance, whereas source-based document filtering leads to a decline in QA performance across all datasets. For instance, on the TriviaQA dataset, the average EM score drops by over 14%. 2) Both methods can only alleviate the "Spiral of Silence" phenomenon to varying degrees and cannot eliminate it. Figure 16 displays the proportion of documents from different sources within the top-5 retrieval results in each iteration under three filtering setups on the NQ dataset. It is observable that without any filtering strategy, human-generated texts rapidly vanish from the top-5 documents in the initial iterations. The SELF-BLEU value filtering method retains human-generated texts to a small extent; source filtering, on the other hand, maximally filters out LLM-generated texts, especially those produced by GPT-3.5-Turbo, Qwen, and ChatGLM3, with over 30% of human-generated texts remaining in the top-5 by the end of the tenth iteration. However, despite both filtering strategies slowing the disappearance of human texts, the proportion of human-generated content continues to exhibit a declining trend. Figure 17 and Figure 18 demonstrate that compared to the absence of filtering strategies, both filtering methods slow down the polarization speed of top document accuracy in retrieval performance. Overall, we discovered that filtering based on the source of documents and their diversity can, to some extent, slow down the emergence of the Spiral of Silence phenomenon. Source-based filtering has a more pronounced effect in terms of preserving the proportion of human-generated texts and mitigating viewpoint polarization; however, this benefit comes at the expense of the performance of the RAG system. Text filtering based on diversity shows superior performance in maintaining RAG system functionality, but it has a weaker impact on preserving the ratio of human texts and alleviating viewpoint polarization. Despite these findings, neither method can completely eradicate the "Spiral of Silence" effect, indicating the imperative to explore additional solutions. For example, there is a need to investigate retrieval models that can effectively balance between LLM-generated documents and human-generated documents to address this issue.
@@
Figure 17: Correlation between the number of top 5 search results containing the correct answer ("Context Right Num") and the accuracy of responses given by LLMs on the NQ dataset when using Content Filtering. The responses are categorized based on Exact Match (EM) score: EM=1 for correct and EM=0 for incorrect. The overall number of queries that the LLMs answered correctly (EM=1 Total) and incorrectly (EM=0 Total), along with the average retrieval accuracy (Acc@5) are shown by dashed lines. The results are averaged across different LLMs, retrieval and ranking methods.
Figure 18: Correlation between the number of top 5 search results containing the correct answer ("Context Right Num") and the accuracy of responses given by LLMs on the NQ dataset when using Source Filtering. The responses are categorized based on Exact Match (EM) score: EM=1 for correct and EM=0 for incorrect. The overall number of queries that the LLMs answered correctly (EM=1 Total) and incorrectly (EM=0 Total), along with the average retrieval accuracy (Acc@5) are shown by dashed lines. The results are averaged across different LLMs, retrieval and ranking methods.
A.8 Prompts

### v3

@@ A.8 Attempts to Alleviate the “Spiral of Silence"
Experimental Setup: To balance the efficiency and effectiveness of the retrieval system, for each set of search results returned by the system, we post-process to acquire the top 5 qualifying documents that are visible to the LLMs. In the source filtering experiment, we employ the Hello-SimpleAI/chatgpt-qa-detector-roberta [^6] https://github.com/Hello-SimpleAI/chatgpt-comparison-detection model to authenticate the origins of the texts within the search results, aiming to retain the first 5 documents identified as human-generated and supply them as input to the LLM’s context. For the content filtering part of the experiment, we apply a selection process based on computing the 3-gram Self-BLEU scores. The specific procedure is as follows: For the top 5 documents returned for each search query, we initially calculate their Self-BLEU scores; if the score exceeds a predetermined threshold (set at 0.4 for this experiment), we then compute the Self-BLEU scores for all possible combinations of 4 documents and select the minimum value among them. This minimum value indicates the maximum individual document contribution to the Self-BLEU score not included in the calculation. Subsequently, we exclude the document contributing the most to the Self-BLEU score and incorporate the next ranked document into the combination, repeating this filtering process until the combination’s Self-BLEU score meets the preset threshold criteria.
@@
Analysis of Experimental Results: From the experimental results, we can observe that: 1) Both Approaches Yield more Stable Retrieval Outcomes; however, the Source Filtering Method Incurs a Performance Cost: Figure 15 and Figure 16 illustrate the variations in the average retrieval outcomes and QA performance across datasets, before and after the application of two distinct filtering strategies, compared to an unfiltered condition. Observations indicate that by implementing source-based and diversity-based filtering methods, the fluctuation range of the top 5 retrieval results is reduced compared to the non-intervention scenario, suggesting that the filtering mechanisms can bring a more stable retrieval performance for RAG systems. Across the four datasets, the retrieval performance following SELF-BLEU value filtering generally surpasses the unfiltered condition; conversely, the source-based filtering strategy results in an overall performance degradation. This could be attributed to the discriminating model erroneously excluding valid human-generated texts while aiming to eliminate those generated by LLMs. Moreover, in QA tasks, diversity filtering either enhances or maintains QA performance, whereas source-based document filtering leads to a decline in QA performance across all datasets. For instance, on the TriviaQA dataset, the average EM score drops by over 14%. 2) Both Methods can only Alleviate the “Spiral of Silence” Phenomenon to Varying Degrees but Cannot Eliminate it: Figure 17 displays the proportion of documents from different sources within the top 5 retrieval results in each iteration under three filtering setups on the NQ dataset. It is observable that without any filtering strategy, human-generated texts rapidly vanish from the top 5 documents in the initial iterations. The SELF-BLEU value filtering method retains human-generated texts to a small extent; source filtering, on the other hand, maximally filters out LLM-generated texts, especially those produced by GPT-3.5-Turbo, Qwen, and ChatGLM3, with over 30% of human-generated texts remaining in the top 5 by the end of the tenth iteration. However, despite both filtering strategies slowing the disappearance of human texts, the proportion of human-generated content continues to exhibit a declining trend. Figure 18 and Figure 19 demonstrate that compared to the absence of filtering strategies, both filtering methods slow down the polarization speed of top document accuracy in retrieval performance. Overall, we discovered that filtering based on the source of documents and their diversity can, to some extent, slow down the emergence of the “Spiral of Silence” phenomenon. Source-based filtering has a more pronounced effect in terms of preserving the proportion of human-generated texts and mitigating viewpoint polarization; however, this benefit comes at the expense of the performance of the RAG system. Text filtering based on diversity shows superior performance in maintaining RAG system functionality, but it has a weaker impact on preserving the ratio of human texts and alleviating viewpoint polarization. Despite these findings, neither method can completely eradicate the “Spiral of Silence” effect, indicating the imperative to explore additional solutions. For example, there is a need to investigate retrieval models that can effectively balance between LLM-generated documents and human-generated documents to address this issue.
@@
Figure 18: Correlation between the number of top 5 search results containing the correct answer ("Context Right Num") and the accuracy of responses given by LLMs on the NQ dataset when using Content Filtering. The responses are categorized based on Exact Match (EM) score: EM=1 for correct and EM=0 for incorrect. The overall number of queries that the LLMs answered correctly (EM=1 Total) and incorrectly (EM=0 Total), along with the average retrieval accuracy (Acc@5) are shown by dashed lines. The results are averaged across different LLMs, retrieval, and ranking methods.
Figure 19: Correlation between the number of top 5 search results containing the correct answer (“Context Right Num”) and the accuracy of responses given by LLMs on the NQ dataset when using Source Filtering. The responses are categorized based on Exact Match (EM) score: EM=1 for correct and EM=0 for incorrect. The overall number of queries that the LLMs answered correctly (EM=1 Total) and incorrectly (EM=0 Total), along with the average retrieval accuracy (Acc@5) are shown by dashed lines. The results are averaged across different LLMs, retrieval and ranking methods.
A.9 Prompts